# Nested Sampling for ARIMA Model Selection in Astronomical Time-Series Analysis

### v1

@@ 3.2 Sampler Configuration
The sampler is initialized by specifying the model order, prior parameters, the number of live points $n_{\text{live}}$ and a random seed for the run. The process is parallelized in this sampler through the simultaneous deletion and repletion of a batch of $n_{\text{delete}}$ lowest-likelihood points. This parallelization parameter is set to $n_{\text{delete}}=50$. The blackjax nested sampler adopts Slice Sampling (Neal, 2003) as a default algorithm for the inner MCMC kernel of the sampler. The length of the MCMC chain run $k$ for each single replacement of the points is expressed in terms of the number of dimensions $\hat{d}$ of the parameter space. This value is conventionally set to $k=6\hat{d}$ in our work. The convergence criterion is defined, as discussed in subsection 2.2.1, using the threshold value:
@@ 3.3 ARIMA Model Selection and Fitting
In this paper, we focus on evaluating the model log posterior probabilities $\log{P_{i}}$ on a grid of ARIMA models formed by the AR order $p$ and the MA order $q$, and a fixed $d$ order. To ensure that the model log posterior probabilities are comparable across candidate ARIMA specifications, each model must be formulated as a complete generative description of the same observed time series. In particular, any two ARIMA models with the same $p$ and $q$ orders but different $d$ orders should be implemented without any external pre-differencing to the data. In Section 4, we demonstrate model selection across the $d$ order for a simulated ARMA process with a trend. In all other cases, for simplicity, we adopt first order differencing if a trend is apparent. The standard stationarity checks prescribed in the Box-Jenkins methodology are also used. In particular, statistical unit root tests such as the Augmented Dickey Fuller (ADF) (Dickey & Fuller, 1979) and Kwiatkowski–Phillips–Schmidt–Shin (KPSS) (Kwiatkowski et al., 1992) Tests are used to confirm stationarity.
A grid search using $n_{\text{live}}<500$ is first performed for identifying the best models. The results are visualized by plotting a heatmap of the calculated model logarithmic posterior probabilities $\log{P_{i}}$ on the ARIMA grid, similar to Yu (2023, Fig.7). From the heatmap, the model with the highest log posterior probability is picked for a second, high resolution nested sampling run.
Further analysis is performed by dividing the time series data into training and observed windows. The model with the highest log posterior probabilities is fit to the training dataset using a high resolution nested sampling run ($n_{\text{live}}=500$ to $1000$). Weighted posterior samples are obtained from this run which are then used to visualize the model fit and residuals. For this, we use fgivenx (Handley, 2018) - a python package for plotting posterior line plots and predictive posteriors of functions. Using the weighted posterior samples, the posterior predictive forecasts along with their $1\sigma$, $2\sigma$ and $3\sigma$ credible regions are plotted.
@@ 3.4 Model Validation
The grid-search results are validated on the broad bases of residual analysis and forecasting accuracy. Practically, almost all ARIMA models are able to model the training window of any given time-series by making one-step ahead, in-sample predictions. On the other hand, a more accurate metric for model validation is comparing the results of out-of-sample multi-step forecasts.
To confirm that the selected best-fit model has correctly captured the variability of the training dataset, we test the residuals from this fit for normality and autocorrelation. This is done by plotting residual histograms and autocorrelation plots along with performing the Ljung-Box statistical test (Ljung & Box, 1978) for autocorrelation. The multi-step direct forecasts for sunspots number data are also tested for accuracy by evaluating the Mean Square Errors (MSE), Root Mean Square Errors (RMSE) and Mean Absolute Errors (MAE). In order to facilitate comparison with a traditional ARIMA model selection method, the Bayesian Information Criterion (BIC) heatmap is also plotted on the ARIMA grid using the ARIMA model functionality from statsmodels.
@@ 4 Results
This section presents the results of applying the model selection procedure to simulated and real astronomical time series data.
@@ 4.1 Simulated Time Series
It suffices to compare the posteriors of the parameters to their true values for the validation of our method in this case. A high resolution nested sampling run of the ARIMA(2,0,0)(2,0,0) model for the time series was able to produce posteriors samples close to the true values of the parameters (Figure 3).
@@
To demonstrate model selection across different $d$ orders of the ARIMA model, we test our framework on a simulated ARMA($1,1$) process with a linear trend (Figure 4). The data was generated using the equation
@@
Nested sampling runs were initiated on this data for ARIMA($1,d,1$) models, with $d$ ranging from 0 to $4$. The calculated model log posterior probabilities $\log P_{i}$ along with their uncertainties are outlined in Table LABEL:logp_d_table.
@@
The posterior probability is highest for ARIMA($1,1,1$) as expected, since first order of differencing is required to eliminate the linear trend in the data.
4.2 Sunspots Number Data
Sunspots are darker and cooler spots observed on the solar photosphere. The evolution of the number of sunspots during the sunspot cycle is found to be directly associated with the solar activity (Hathaway, 2015). It is therefore essential to analyse and forecast the sunspots number for predicting space weather and mitigating its impact on Earth.
We use the yearly sunspots number data from the year of 1700 to 2008 for our analysis.
Figure 5: Yearly Sunspots Number Data from 1700 to 2008
@@ 4.2 Sunspots Number Data
Table 3: Results of the Stationarity Tests (ADF and KPSS) on yearly sunspots number data.
@@
From the p-value and the critical value at 5% significance level of both tests, it can be concluded that the time-series data is stationary. Therefore, no differencing is required and the $d$ order is set to zero for the ARIMA grid search. A grid search was carried out for ARIMA orders up-to ten (Figure 6), which selected ARIMA($9,0,1$) model with the highest log posterior probability of $\log{P_{\text{max}}}=-1.3\pm 0.5$.
@@
Figure 6: Heatmap of the model log posterior probabilities obtained from the nested sampling runs on yearly sunspots number data (Figure 5). The grid is annotated with model log posterior probabilities $\log{P_{i}}$ and their estimated uncertainties $\sigma_{\log{P_{i}}}$ for reference. A clear statistical preference for higher ARIMA orders, and particularly for a higher AR order $p$ can be seen, with ARIMA(9,0,1)(9,0,1) having the highest model posterior probability. .
On the contrary, ARIMA($3,0,3$) is chosen as the best-fit model on the basis of the Bayesian Information Criterion (BIC), as shown in Figure 7. The maximum likelihood estimation method adopted by the ARIMA model function in statsmodels failed to converge for multiple higher order ARIMA models.
Figure 7: Heatmap of the Bayesian Information Criterion (BIC) values for ARIMA model fit on yearly sunspots number data. The lowest value of BIC is observed for ARIMA($3,0,3$). There is no clear preference seen for higher ARIMA orders suggesting that this method of model selection has penalized complex models more strongly than nested sampling.
@@
ARIMA(9,0,1)(9,0,1) is chosen for fitting the data using a high resolution ($n_{\text{live}}=1000$) nested sampling run. The marginalized posterior distributions of ARIMA parameters obtained from this fit are shown in Figure 8.
@@
Figure 8: Posterior corner plot of ARIMA($9,0,1$) model parameters obtained from the nested sampling run on the yearly sunspots number data (Figure 5). Orange curves represent the marginalized posterior distributions, while the gray curves show the corresponding prior distributions (Table 1) adopted for each parameter. The sharp posterior peaks indicate well-constrained AR and MA coefficients, with most of the posterior mass concentrated within the stationary and invertible regions of parameter space. The information gain between prior and posterior, measured by the Kullback–Leibler divergence, is $D_{KL}=19$ nats, demonstrating a substantial concentration of probability mass due to the data.
@@
Figure 9: Results of ARIMA(9,0,1)(9,0,1) model fit to the training window, corresponding to the years of from 1700 to 1954, of the yearly sunspots number data (Figure 5). The upper panel shows the observed sunspot counts (in black) and the model fit curves (in red) obtained using $500$ weighted posterior samples of the ARIMA parameters, while the lower panel displays the corresponding residuals. The residuals fluctuate around zero with no strong temporal structure, indicating that the fitted model captures most of the systematic variability in the data.
The model occasionally predicts a negative sunspots number count due to the assumption of normally distributed errors $\epsilon_{t}\sim\mathcal{N}(0,\sigma^{\prime})$. We regard these predictions as unphysical. It can be seen that the residual sequence symmetrically fluctuates around zero. A histogram of the pooled residuals from different posterior samples is shown in Figure 10, depicting the residuals to be almost normally distributed.
Figure 10: A histogram of the pooled residuals from the ARIMA(9,0,1)(9,0,1) fit in Figure 9 to yearly sunspots number data. The residuals appear to be normally distributed around 0.
@@
Figure 11: Autocorrelation (top) and Partial Autocorrelation (bottom) function plots of the mean residuals from the ARIMA(9,0,1)(9,0,1) fit to the yearly sunspots number data (Figure 9). Both functions lie within the 95% confidence bounds across all lags, indicating that the residuals are effectively uncorrelated.
@@
Table 4: Ljung–Box(LB) test p-values for residual autocorrelation of sunspots number data.
@@
The results show that all p-values exceed $0.05$ across lags 11-$10$. Hence, we fail to reject the null hypothesis of no serial correlation in the residuals at the $5\%$ significance level. The residuals are therefore characteristic of a white noise sequence, indicating that the ARIMA(9,0,1)(9,0,1) fit have effectively captured the temporal structure of our training data.
Finally, we show in Figure 12, the results of direct multi-step forecasting of the sunspots number data from the years of 1954 to 2008, using an ARIMA(9,0,1)(9,0,1) fit.
Figure 12: Posterior predictive forecasts of the yearly sunspots number for the years from 1954 to 2008. The forecasts are obtained using $5000$ weighted posterior samples from the ARIMA(9,0,1)(9,0,1) fit (Figure 9). Shaded contours denote the $1\sigma$, $2\sigma$ and $3\sigma$ credible regions of the predictive posterior distribution $P(\hat{y}_{t}|t,D_{t})$.
@@ 4.3 Kepler and TESS lightcurves
The Kepler and TESS (Transiting Exoplanet Survey Satellite) photometric lightcurves act as ideal datasets for testing our ARIMA-Nested Sampling framework owing to their long and regular cadences with minimal systematics noise. In this subsection, we present the results of applying our model selection methodology to photometric lightcurves of two targets observed by Kepler and TESS - KIC 12008916 and Ross 176, respectively.
@@
KIC 12008916 is a low-luminosity red giant in the original Kepler field and is one of the benchmark stars in the Kepler astroseismic sample (Davies & Miglio, 2016). The star has a well-resolved, solar-like oscillation spectrum. The lightcurve exhibits stochastic variability which is stable and stationary over longer timescales, providing an ideal testbed for ARIMA modelling.
@@
The Pre-search Data Conditioning Search Aperture Photometry flux (PDCSAP) flux was used for this analysis, which was found to be significantly autocorrelated, as seen from the ACF/PACF plots in the top panel of Figure 15. The results of the ADF and KPSS stationarity tests confirm the time series data to be stationary. An ARIMA grid search, with $d=0$ fixed, revealed ARIMA(0,0,1)(0,0,1) as the best fit model ($\log{P_{\text{max}}}=-0.0014\pm 0.76$) despite the significant autocorrelation present beyond first lag in the data. The results of the grid search are shown in Figure 14.
@@
Figure 14: Heatmap of the ARIMA models’ log posterior probabilities obtained from the nested sampling runs on KIC 12008916 data (Figure 13).
@@
The ARIMA(0,0,1)(0,0,1) model was fit to the data using a high resolution ($n_{\text{live}}=500$) nested sampling run and the means of the posterior samples were used to obtain time-ordered residuals. We observe no significant autocorrelation in these time-ordered residuals, as shown by their ACF/PACF plots from the bottom panel of Figure 15.

### v2

@@ 3.2 Sampler Configuration
The sampler is initialized by specifying the model order, prior parameters, the number of live points $n_{\text{live}}$ and a random seed for the run. The process is parallelized in this sampler through the simultaneous deletion and replacement of a batch of $n_{\text{delete}}$ lowest-likelihood points. This parallelization parameter is set to $n_{\text{delete}}=50$. The blackjax nested sampler adopts Slice Sampling (Neal, 2003) as a default algorithm for the inner MCMC kernel of the sampler. The length of the MCMC chain run $k$ for each single replacement of the points is expressed in terms of the number of dimensions $\hat{d}$ of the parameter space. This value is conventionally set to $k=6\hat{d}$ in our work. The convergence criterion is defined, as discussed in subsection 2.2.1, using the threshold value:
@@ 3.3 ARIMA Model Selection and Fitting
In this paper, we focus on evaluating the model log posterior probabilities $\log{P_{i}}$ on a grid of ARIMA models formed by the AR order $p$ and the MA order $q$, and a fixed $d$ order. To ensure that the model log posterior probabilities are comparable across candidate ARIMA specifications, each model must be formulated as a complete generative description of the same observed time series. In particular, any two ARIMA models with the same $p$ and $q$ orders but different $d$ orders should be implemented without any external pre-differencing to the data. In Section 4, we demonstrate model selection across the $d$ order for a simulated ARMA process with a trend. In all other cases, for simplicity, we fix a priori the order of differencing $d$ using standard stationarity checks on the time series as prescribed in the Box-Jenkins methodology. In particular, statistical unit root tests such as the Augmented Dickey-Fuller (ADF) (Dickey & Fuller, 1979) and Kwiatkowski–Phillips–Schmidt–Shin (KPSS) (Kwiatkowski et al., 1992) Tests are used to confirm stationarity.
A grid search using $n_{\text{live}}=100$ is first performed to identify the best models. The results are visualized by plotting a heatmap of the calculated model logarithmic posterior probabilities $\log{P_{i}}$ on the ARIMA grid, similar to Yu (2023, Fig.7). From the heatmap, the model with the highest log posterior probability is picked for a second, high-resolution nested sampling run.
Further analysis is performed by dividing the time series data into training and test windows. The model with the highest log posterior probabilities is fit to the training dataset using a high-resolution nested sampling run ($n_{\text{live}}=500$ to $1000$). Weighted posterior samples are obtained from this run, which are then used to visualize the model fit and residuals. For this, we use fgivenx (Handley, 2018) - a python package for plotting posterior line plots and predictive posteriors of functions. Using the weighted posterior samples, the posterior predictive forecasts along with their $1\sigma$, $2\sigma$ and $3\sigma$ credible regions are plotted.
@@ 3.4 Model Validation
The grid-search results are validated on the basis of residual analysis and forecasting accuracy. Practically, almost all ARIMA models are able to model the training window of any given time-series by making one-step-ahead, in-sample predictions. On the other hand, a more accurate metric for model validation is comparing the results of out-of-sample multi-step forecasts.
To confirm that the selected best-fit model has correctly captured the variability of the training dataset, we test the residuals from this fit for normality and autocorrelation. This is done by plotting residual histograms and autocorrelation plots along with performing the Ljung-Box statistical test (Ljung & Box, 1978) for autocorrelation. The multi-step direct forecasts for sunspot number data are also tested for accuracy by evaluating the Mean Squared Error (MSE), Root Mean Squared Error (RMSE) and Mean Absolute Error (MAE). In order to facilitate comparison with a traditional ARIMA model selection method, the Bayesian Information Criterion (BIC) heatmap is also plotted on the ARIMA grid using the ARIMA model functionality from statsmodels.
For time series other than the sunspot number, our interest is not in forecasting performance but in the model’s ability to accurately capture the variability and correlation in the given time series. Therefore, a simple residual analysis, as described previously, suffices in such cases.
@@ 4 Results
This section presents the results of applying the model selection procedure to simulated and real astronomical time series data using the previously described methodologies.
@@ 4.1 Simulated Time Series
Comparing the inferred posteriors to the true parameter values suffices to validate the method in this case. A high-resolution nested sampling run of the ARIMA(2,0,0)(2,0,0) model for the time series was able to produce posterior samples close to the true values of the parameters (Figure 3).
@@
To demonstrate model selection across different $d$ orders of the ARIMA model, we test our framework on a simulated ARMA($1,1$) process with a linear trend (Figure 4). The data were generated using the equation
@@
Nested sampling runs were initiated on these data for ARIMA($1,d,1$) models, with $d$ ranging from 0 to $4$. The calculated model log posterior probabilities $\log P_{i}$ along with their uncertainties are outlined in Table LABEL:logp_d_table.
@@
The posterior probability is highest for ARIMA($1,1,1$) as expected, since first-order differencing is required to eliminate the linear trend in the data.
4.2 Sunspot Number Data
Sunspots are darker and cooler spots observed on the solar photosphere. The evolution of the number of sunspots during the sunspot cycle is found to be directly associated with the solar activity (Hathaway, 2015). It is therefore essential to analyse and forecast the sunspot number for predicting space weather and mitigating its impact on Earth.
We use the yearly sunspot number data from 1700 to 2008 for our analysis.
Figure 5: Yearly sunspot number data from 1700 to 2008
@@ 4.2 Sunspot Number Data
Table 3: Results of the Stationarity Tests (ADF and KPSS) on yearly sunspot number data.
@@
From the p-value and the critical value at 5% significance level of both tests, it can be concluded that the time-series data is stationary. Therefore, no differencing is required and the $d$ order is set to zero for the ARIMA grid search. A grid search was carried out for ARIMA orders up to ten (Figure 6), which selected the ARIMA($9,0,1$) model with the highest log posterior probability of $\log{P_{\text{max}}}=-1.3\pm 0.5$.
@@
Figure 6: Heatmap of the model log posterior probabilities obtained from the nested sampling runs on yearly sunspot number data (Figure 5). The grid is annotated with model log posterior probabilities $\log{P_{i}}$ and their estimated uncertainties $\sigma_{\log{P_{i}}}$ for reference. A clear statistical preference for higher ARIMA orders, and particularly for a higher AR order $p$ can be seen, with ARIMA(9,0,1)(9,0,1) having the highest model posterior probability.
The maximum likelihood estimation method adopted by the ARIMA model function in statsmodels failed to converge for multiple higher-order ARIMA models (see Figure 7). ARIMA(3,0,3)(3,0,3) was picked as the best-fit model using this method by minimizing the Bayesian Information Criterion (BIC). To confirm this is not an artefact of poor initialisation, we additionally ran auto_arima from the pmdarima library with a full grid search up to $p=q=10$ and BIC as the selection criterion. This returned ARIMA(3,0,2)(3,0,2) as the best-fit model and also reported convergence failures at some higher-order models, consistent with the statsmodels result. The close model preference between two independent MLE implementations suggests that this disagreement with our nested sampling result reflects a fundamental limitation of likelihood optimisation in high-dimensional ARMA parameter spaces rather than a calibration issue. As demonstrated in Appendix A, ARIMA(9,0,1)(9,0,1) selected by nested sampling achieves substantially better performance metrics on the held-out test window than ARIMA(3,0,3)(3,0,3), providing empirical validation that the Bayesian evidence identifies a richer model structure than MLE-based methods.
Figure 7: Heatmap of the Bayesian Information Criterion (BIC) values for ARIMA model fit on yearly sunspot number data. Missing data indicate models for which the maximum likelihood estimation failed to converge. The lowest value of BIC is observed for ARIMA($3,0,3$)
The preference for a high AR order of $p=9$ in the heatmap data (Figure 6) has a physically motivated interpretation. Solar magnetic activity is known to operate across a hierarchy of timescales. The dominant periodicity is the Schwabe cycle of approximately 11 years, but significant structure also exists on shorter timescales associated with active region emergence, differential rotation, and the stochastic nature of flux tube emergence at the solar surface. In an AR model, the order $p$ determines the memory depth of the process in units of the sampling interval; with annual sampling, an AR order of 9 therefore captures dependencies extending approximately 9 years into the past, commensurate with the rise phase of the solar cycle. The single MA term ($q=1$) is consistent with a short-memory noise process superimposed on this long-memory autoregressive structure.
@@
We choose ARIMA(9,0,1)(9,0,1) for fitting the data using a high-resolution ($n_{\text{live}}=1000$) nested sampling run. The marginalized posterior distributions of ARIMA parameters obtained from this fit are shown in Figure 8.
@@
Figure 8: Posterior corner plot of ARIMA($9,0,1$) model parameters obtained from the nested sampling run on the yearly sunspot number data (Figure 5). Orange curves represent the marginalized posterior distributions, while the gray curves show the corresponding prior distributions (Table 1) adopted for each parameter. The sharp posterior peaks indicate well-constrained AR and MA coefficients, with most of the posterior mass concentrated within the stationary and invertible regions of parameter space. The information gain between prior and posterior, measured by the Kullback–Leibler divergence, is $D_{KL}=19$ nats, demonstrating a substantial concentration of probability mass due to the data.
@@
Figure 9: Results of ARIMA(9,0,1)(9,0,1) model fit to the training window, corresponding to the years of from 1700 to 1954, of the yearly sunspot number data (Figure 5). The upper panel shows the observed sunspot counts (in black) and the model fit curves (in red) obtained using $500$ weighted posterior samples of the ARIMA parameters, while the lower panel displays the corresponding residuals. The residuals fluctuate around zero with no strong temporal structure, indicating that the fitted model captures most of the systematic variability in the data.
We note that the model occasionally produces negative predicted sunspot counts, which are unphysical. This is a known limitation of applying a Gaussian likelihood to strictly non-negative count data. A more robust treatment would adopt a truncated Gaussian or a negative binomial likelihood. It can be seen that the residual sequence symmetrically fluctuates around zero. A histogram of the pooled residuals from different posterior samples is shown in Figure 10, depicting the residuals to be almost normally distributed.
Figure 10: A histogram of the pooled residuals from the ARIMA(9,0,1)(9,0,1) fit in Figure 9 to yearly sunspot number data. The residuals appear to be normally distributed around 0.
@@
Figure 11: Autocorrelation (top) and Partial Autocorrelation (bottom) function plots of the mean residuals from the ARIMA(9,0,1)(9,0,1) fit to the yearly sunspot number data (Figure 9). Both functions lie within the 95% confidence bounds across all lags, indicating that the residuals are effectively uncorrelated.
@@
Table 4: Ljung–Box(LB) test p-values for residual autocorrelation of sunspot number data.
@@
The results show that all p-values exceed $0.05$ across lags 11-$10$. Hence, we fail to reject the null hypothesis of no serial correlation in the residuals at the $5\%$ significance level. The residuals are therefore characteristic of a white noise sequence, indicating that the ARIMA(9,0,1)(9,0,1) fit has effectively captured the temporal structure of our training data.
Finally, we show in Figure 12, the results of direct multi-step forecasting of the sunspot number data from 1954 to 2008, using an ARIMA(9,0,1)(9,0,1) fit.
Figure 12: Posterior predictive forecasts of the yearly sunspot number for the years from 1954 to 2008. The forecasts are obtained using $5000$ weighted posterior samples from the ARIMA(9,0,1)(9,0,1) fit (Figure 9). Shaded contours denote the $1\sigma$, $2\sigma$ and $3\sigma$ credible regions of the predictive posterior distribution $P(\hat{y}_{t}|t,D_{t})$.
@@ 4.3 Kepler and TESS light curves
The Kepler and TESS (Transiting Exoplanet Survey Satellite) photometric light curves act as ideal datasets for testing our ARIMA-Nested Sampling framework owing to their long and regular cadences with minimal systematics noise. In this subsection, we present the results of applying our model selection methodology to photometric light curves of Kepler and TESS. We focus on analyzing the light curves of three distinct astrophysical systems: a low-luminosity red giant – KIC 12008916, an active G-type star hosting a hot Jupiter – Kepler 17, and two quasars observed by TESS – 3C 273 and S4 0954+65.
@@
KIC 12008916 is a low-luminosity red giant in the original Kepler field and is one of the benchmark stars in the Kepler asteroseismic sample (Davies & Miglio, 2016). The star has a well-resolved, solar-like oscillation spectrum. Its light curve exhibits stochastic variability which is stable and stationary over longer timescales, providing an ideal test-bed for ARIMA modelling.
@@
The Pre-search Data Conditioning Simple Aperture Photometry (PDCSAP) flux was used for this analysis, which was found to be significantly autocorrelated, as seen from the Autocorrelation Function (ACF) and Partial Autocorrelation Function (PACF) plots in the top panel of Figure 17. The results of the ADF and KPSS stationarity tests confirmed the time series data to be stationary.
Table 5: Results of Augmented Dickey-Fuller (ADF) and KPSS Stationarity Tests for KIC 12008916 light curve.
Statistic ADF Test KPSS Test
Test Statistic -4.620294 0.100269
p-value 0.000118 0.100000
Lags Used 9 10
Number of Observations Used 459
Critical Value (1%) -3.444677 0.739000
Critical Value (5%) -2.867857 0.463000
Critical Value (10%) -2.570135 0.347000
An ARIMA grid search, with $d=0$ fixed, revealed ARIMA(0,0,4)(0,0,4) as the best-fit model ($\log{P_{\text{max}}}=-0.18\pm 0.67$). The results of the grid search are shown in Figure 14.
@@
Figure 14: Heatmap of the ARIMA models’ log posterior probabilities obtained from the nested sampling runs on KIC 12008916 data (Figure 13). A statistical preference for purely moving average models is seen, with the highest log posterior probability for ARIMA(0,0,4)(0,0,4).
@@
The ARIMA(0,0,4)(0,0,4) model was fit to the data using a high-resolution ($n_{\text{live}}=500$) nested sampling run and the means of the posterior samples were used to obtain time-ordered residuals (see Figure 15).
Figure 15: Results of ARIMA(0,0,4)(0,0,4) model fit to the KIC 12008916 light curve in Figure 13. Top panel shows the light curve data (in black) and the model fit curve (in red) obtained from the mean of the posterior samples. Bottom panel shows the residuals obtained after subtracting the mean model fit from the data. The residual time series show no temporal structure and is characteristic of a white noise sequence.
A histogram of these residuals is shown in Figure 16. The residuals appear to be normally distributed around zero.
Figure 16: A histogram of the residuals in Figure 15, obtained from the ARIMA(0,0,4)(0,0,4) model fit to the light curve data of KIC 12008916. The residuals appear to be normally distributed around 0.
The autocorrelation and partial autocorrelation plots of these residuals are also plotted in Figure 17. We observe no significant autocorrelation in the time-ordered residuals, as seen from these plots.